OpenStudy (anonymous):

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (anonymous):

Lets start with part c he made $490 in a week Divide that by ten then show me what u get.

OpenStudy (anonymous):

49

OpenStudy (anonymous):

yep now he can only work 40 hours with making ten dollars. so we know that he worked 40 hours for sure. now for part a they want to know how much money he made earning 10 dollars an hour without going over time. do what is 40 * 10 = ?

OpenStudy (anonymous):

400

OpenStudy (anonymous):

Great Job! so our equation for part A would be $10 * 40 = $400 correct?

OpenStudy (anonymous):

do you agree^^^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

making sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok awesome! now for part B they want you to create an equation for the amount he worked overtime. Now Do you remember how many hours a week he can work without going over time?

OpenStudy (anonymous):

any hours more than 40 hours a week

OpenStudy (anonymous):

Correct! and we already figured out he worked 49 hours this week. so that means he worked 9 hour overtime. and he makes $15 an hour over time so what is 15*9=?

OpenStudy (anonymous):

understand?